andrew earns an average of 90 in homework and 80 in quizzes. What final examination grade will give him a total average of 83 if homework has a weight of 1/6 , quizzes 3/6 and final examination 2/6??
Answer by jorel555(1290)

If andrew averages 90 in homework, and 80 in quizzes, which count for 1/6 and 1/2 of his total grade, then:
90/6+80/2=15+40=55
83-55=28
Since the final exam counts as 1/3 of his grade, then he needs to score 28/(1/3)=28 x 3= 84 on his final exam to obtain an 83 average.